OpenStudy (anonymous):

Could you give me the prime factorizations? 64x^2 -176xy + 121y^2

OpenStudy (anonymous):

a^2 -2ab + b^2 where a = 8x b = 11y

OpenStudy (anonymous):

can you show how to factor this complety?

OpenStudy (anonymous):

\[(8x)^2 - 2\times{8x}\times{11y} + (11y)^2\] \[(8x-11y)^2\]

OpenStudy (anonymous):

Megan factored the expression as72x^2-43x+6 as (9x-3)(8x-2) . But when Jacob applied the FOIL principle and multiplied out (9x-3)(8x-2), he got 72x^2-42x+6 ; thus, Megan’s solution does not appear to check. Why is that? Please help Megan to understand this better. Explain your reasoning and correctly factor the original expression, if possible. If the expression is prime, so state.